OpenStudy (anonymous):

The perimeter of a recyangle is 84cm. The length of the rectangle is six more than the width. Find the length

OpenStudy (ghazi):

perimeter = 2(length + breadth), just substitute 84= 2(6*breadth+breadth), you'll get breadth = 6 hence find length

OpenStudy (anonymous):

Choices ate 30 27 24 18

OpenStudy (ghazi):

length = 6*breadth....therefore length= 36 i guess your options are wrong...

OpenStudy (anonymous):

How did u het 36?

OpenStudy (ghazi):

perimeter = 2*(length + breadth) and it's given length= 6*breadth therefore substitute breadth in the equation ..you'll get breadth= 6 and length =6*breadth=6*6=36

OpenStudy (anonymous):

Thanks

OpenStudy (ghazi):

anytime